OpenStudy (anonymous):

Confused on what to do... Find the product:

OpenStudy (anonymous):

\[x ^{4}x ^{3}\]

OpenStudy (anonymous):

x^7 is the answer bud

OpenStudy (anonymous):

Don't you just add the 4+3? so it becomes x^7?

OpenStudy (anonymous):

Oh okay haha(: Thanks!

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <--- iuse the 1s rule

OpenStudy (anonymous):

you are just multiplying the two variables. When multiplying variables with exponents, add the exponents. Thus, 4+3=7

OpenStudy (anonymous):

Then what would I do for \[w ^{8}w\]

OpenStudy (anonymous):

w^9

OpenStudy (anonymous):

8+1=9. The w without an exponent showing really is w^1

OpenStudy (anonymous):

Ohhh! Yeah I remember that now!(:

OpenStudy (anonymous):

The addition of exponents applies whenever multiplying variables

OpenStudy (anonymous):

\[9x ^{5}(4x ^{7})\]

OpenStudy (anonymous):

Come on, I think you can figure it out now. What is 9x4=? What does 5+7=?

OpenStudy (anonymous):

36^12?

OpenStudy (anonymous):

36x^12. Dont forget the variable

OpenStudy (anonymous):

Okay !(: So I just multiply variables and add exponents!

OpenStudy (anonymous):

Multiply any number in front of the variables, add the exponents :)